Biology pH Calculation

A solution of OH ions is added to a beaker of pure water at 25°C, to a final concentration of 688 mM.

Calculate the pH of the solution
Answer #1

pOH= - log (OH-)  

pH= - Log (H+)

pH+ pOH= 10^14

Please look in the image for pH of the solution: